### v1

# Thought2Text: Text Generation from EEG Signal using Large Language Models (LLMs)

Abstract
Decoding and expressing brain activity in a comprehensible form is a challenging frontier in AI. This paper presents Thought2Text, which uses instruction-tuned Large Language Models (LLMs) fine-tuned with EEG data to achieve this goal. The approach involves three stages: (1) training an EEG encoder for visual feature extraction, (2) fine-tuning LLMs on image and text data, enabling multimodal description generation, and (3) further fine-tuning on EEG embeddings to generate text directly from EEG during inference. Experiments on a public EEG dataset collected for six subjects with image stimuli demonstrate the efficacy of multimodal LLMs (LLaMa-v3, Mistral-v0.3, Qwen2.5), validated using traditional language generation evaluation metrics, GPT-4 based assessments, and evaluations by human expert. This approach marks a significant advancement towards portable, low-cost "thoughts-to-text" technology with potential applications in both neuroscience and natural language processing (NLP).
1 Introduction
Brain-Computer Interface (BCI) systems, combined with portable and wearable noninvasive Electroencephalographic (EEG) devices, enable direct interfacing between the brain and external devices He et al. (2015). Advances in generating images and natural language using EEG Speier et al. (2016); Benchetrit et al. (2023); Défossez et al. (2023) hold promise for developing BCIs in various domains, including assistive communication (e.g., for ALS and stroke patients), mixed reality (AR/VR) experience enhancement, mental health diagnosis, and gaming. Recent strides in NLP driven by powerful LLMs such as OpenAI GPT-4. Achiam et al. (2023), Google Gemini Team et al. (2023), Meta-LLaMa Touvron et al. (2023), Mistral Jiang et al. (2023), and Microsoft Phi Gunasekar et al. (2023) have enabled multimodal integration, facilitating language generation from images Liu et al. (2024b) and speech Fathullah et al. (2024). Our research focuses on a multimodal solution to decode brain signals directly into text, leveraging EEG signals. We choose EEG for its affordability as compared to mainstream alternatives like functional Magnetic Resonance Imaging (fMRI) Tang et al. (2023), which are costlier and require complex setup.
Our approach for generating textual descriptions from EEG signals involves three key steps: (a) capturing language-agnostic EEG signals via visual stimuli, (b) encoding these signals into embeddings using a deep multichannel neural encoder, and (c) fine-tuning language models by projecting image and EEG embeddings into a token embedding space to generate responses. These responses are compared with gold standard image descriptions to compute the training loss. During inference, only EEG signals and a generic textual prompt are used as inputs to the LLMs to generate responses. Our method requires images, EEG data, and descriptions for training, while inference is bimodal, using only EEG to generate text.
For experiments, we used a public 128-channel EEG dataset from six participants viewing visual stimuli. Image descriptions were generated by GPT-4-Omni Achiam et al. (2023) and validated by experts, providing the necessary text modality for building EEG-to-text generation systems. Although the goal is text generation from EEG, we use a dataset with visual stimuli for their language-agnostic nature, avoiding cognitive biases linked to text reading (see Section 3 for details). We fine-tuned large language models (LLMs) using these descriptions, leveraging pre-trained instruction-based language models such as Mistral-v3 Jiang et al. (2023), LLaMa-v3 Touvron et al. (2023), and Qwen2.5 Bai et al. (2023a). Evaluation with standard generation metrics Sharma et al. (2017) and GPT-4-based assessments confirmed our approach’s effectiveness.
Our paper’s key contributions include:
Integration of brain signals with LLMs.
Fine-tuning models on EEG signals captured for visual stimuli, leveraging its language-agnostic nature to enhance LLM interaction and scalability across languages.
Validation of model efficacy on a popular public dataset that contains EEG signals captured using affordable devices in less constrained environments compared to traditional methods.
The code and a link to the processed dataset can be found at https://github.com/abhijitmishra/Thought2Text.
2 Related Work
Integrating behavioral signals like eye-gaze and brain signals into NLP and computer vision tasks Mishra and Bhattacharyya (2018); Sharma and Meena (2024) has seen significant progress. Key datasets include ZuCo 2.0 Hollenstein et al. (2019), which captures EEG and eye-gaze during natural language reading, and MOABB Jayaram and Barachant (2018), offering over 120,000 EEG samples from 400+ subjects from various BCI tasks such as motor imagery, visual evoked potentials, and cognitive load. Datasets such as CVPR2017 Spampinato et al. (2017) and MindBigData Vivancos and Cuesta (2022) provide substantial EEG data from responses to handwritten and open vocabulary object-based image stimuli. These datasets have facilitated research in classifying EEG data Spampinato et al. (2017); Palazzo et al. (2020); Khaleghi et al. (2023) and generating images from EEG signals using GANs Goodfellow et al. (2020) and latent diffusion models Rombach et al. (2022); Bai et al. (2023b); Lan et al. (2023); Tirupattur et al. (2018). However, using EEG signals in language generation remains elusive. The closest work to ours is EEG2TEXT Liu et al. (2024a), which uses EEG pretraining and a multi-view transformer to decode EEG signals into text. Unlike our approach, it does not leverage LLMs with prompt engineering for prompt-specific responses. We believe our method of fine-tuning LLMs with EEG input is the first of its kind.
| Annotator | Percentage of Correct Captions |
| Annotator 1 | 98% |
| Annotator 2 | 96% |
| Agreement Score | 93% |
3 Dataset and the Need for Visual Stimuli
Building a system that generates text from thoughts naturally requires a dataset of paired examples. However, we believe that directly presenting participants with textual stimuli is not ideal, as using EEG data collected in this manner for thought analysis introduces additional complexities of language processing, such as determining brain activity windows for specific words, handling the retention of word context post-onset, and managing the overlap of contexts when words are shown in different time frames Wehbe et al. (2014); Murphy et al. (2022). Additionally, vocabulary size presents a challenge: while EEG-to-text systems perform well in closed-vocabulary settings, open-vocabulary decoding becomes inefficient as vocabulary size increases Martin et al. (2018); Wang and Ji (2022); Liu et al. (2024a). The core challenge in thought-to-text systems thus lies in two key aspects: (a) harnessing language-agnostic neural signals with minimal interference from linguistic processing and (b) using these signals to generate text in a target language, potentially by specifying an instruction or prompt.
This leads us to favor datasets where EEG signals are recorded in response to visual stimuli rather than textual input. For (a), visual stimuli provide several advantages: by relying on images, it circumvents the complexities of language processing and also elicits brain responses to salient image features, making them more suited for capturing neural activity in a language-agnostic manner. For (b), generating text from visual stimuli-evoked EEG remains challenging, as most EEG datasets collected with visual stimuli (e.g., CVPR2017 dataset Palazzo et al. (2020), MOABB Jayaram and Barachant (2018), MindBigData Vivancos and Cuesta (2022)) lack an associated text component, which is essential for evaluating whether the generated text from EEG signals accurately captures the perceived salient features of the images. To address this, robust image captioning tools like GPT-4 Omni Achiam et al. (2023) can be employed to generate captions for the visual stimuli, thereby creating a trimodal dataset with pairs. The text generation can further be guided by specifying a prompt or instruction to tailor the output to a particular language or context. This approach enriches the dataset and provides a more comprehensive foundation for thought-to-text research.
For our experiments, we utilize the CVPR2017 dataset Palazzo et al. (2020), which contains EEG data from six participants, each viewing 50 images across 40 diverse object categories (such as vehicles, musical instruments, etc.). Each EEG recording, corresponding to one participant and one visual stimulus, consists of 128 channels recorded for 0.5 seconds at a sampling rate of 1 kHz. This data is represented as a 128×N matrix, where N, approximately 500, represents the number of samples per channel in each segment. Since the exact duration of the EEG signals may vary, we follow the preprocessing approach of Palazzo et al. (2020) by discarding the first 20 samples (20 ms) to reduce interference from the previous image, and then standardizing the signal to a common length of 440 samples, accounting for segments with N < 500. The dataset provides pre-filtered signals across three frequency ranges: 14-70Hz, 5-95Hz, and 55-95Hz. In line with prior research, we select the 55-95 Hz range, as it has demonstrated the most reliable results.
Since the CVPR2017 dataset lacks textual descriptions, we generate one-line captions for each image using GPT-4. We choose brief captions because the half-second EEG recordings may not capture enough information to support longer, more detailed descriptions. Given that these captions are auto-generated, it is crucial to ensure their quality and relevance. To validate the accuracy of the captions, we employed human annotators via Amazon Mechanical Turk (https://www.mturk.com/). These human experts reviewed the images and assessed the captions on a binary scale, judging both their fluency and adequacy. Table 1 summarizes the validation results, including the level of agreement between annotators when both rated a caption as acceptable, demonstrating the reliability of the auto-generated captions. We will release the descriptions under the same license as the original data.
| LLM | ROUGE-N | ROUGE-L | BLEU-N | METEOR | BERT | GPT-4 | GPT-4 | ||
| N=1 | N=2 | N=1 | N=4 | Score | Fluency | Adequacy | |||
| 9.8 | 1.5 | 8.5 | 7.3 | 1.3 | 12.6 | 0.84 | 3.44 | 1.30 | |
| 3.8 | 0.4 | 3.3 | 2.8 | 0.4 | 5.9 | 0.84 | 4.72 | 1.08 | |
| 28.9 | 7.3 | 26.2 | 24.1 | 5.2 | 23.7 | 0.89 | 4.80 | 1.49 | |
| 26.9 | 6.1 | 23.9 | 22.6 | 4.3 | 23.7 | 0.88 | 4.83 | 1.41 | |
| 30.0 | 8.1 | 26.6 | 25.5 | 5.5 | 26.3 | 0.89 | 4.82 | 1.58 | |
| 17.6 | 3.4 | 14.8 | 14.5 | 2.5 | 23.2 | 0.86 | 4.46 | 1.52 | |
| 17.9 | 3.6 | 15.1 | 15.7 | 2.9 | 22.8 | 0.87 | 4.89 | 1.55 | |
| 26.7 | 5.3 | 23.5 | 23.3 | 4.2 | 22.0 | 0.88 | 4.82 | 1.25 | |
| 29.2 | 7.3 | 26.5 | 24.1 | 5.0 | 24.0 | 0.89 | 4.77 | 1.60 | |
| 30.6 | 8.8 | 28.0 | 26.0 | 6.1 | 26.2 | 0.89 | 4.79 | 1.65 | |
| 17.6 | 2.8 | 14.5 | 14.8 | 2.4 | 21.0 | 0.85 | 3.91 | 1.47 | |
| 1.7 | 0.1 | 1.6 | 1.3 | 0.3 | 6.4 | 0.84 | 4.73 | 1.01 | |
| 25.2 | 3.6 | 21.5 | 21.9 | 3.2 | 20.2 | 0.88 | 4.77 | 1.10 | |
| 24.4 | 4.1 | 20.9 | 20.7 | 3.3 | 20.2 | 0.88 | 4.66 | 1.24 | |
| 26.4 | 4.6 | 22.8 | 22.7 | 3.7 | 21.1 | 0.88 | 4.75 | 1.28 | |
4 Method
Our approach employs a three-stage process to generate coherent text from EEG signals. The overall workflow is illustrated in Figure 1.
4.1 Stage1: Training EEG Encoder for Embedding Extraction
The first stage of our approach focuses on developing an encoder that extracts meaningful embeddings () from multi-channel EEG signals. Since the target thoughts are short and pertain to the most salient features of an image, we design the encoder with two objectives: (a) aligning the EEG embeddings with those derived from image stimuli using a pretrained visual encoder, and (b) predicting the most salient object (e.g., piano) from the EEG embeddings.
As shown in Figure 1, we utilize a multichannel EEG encoder inspired by ChannelNet Palazzo et al. (2020), a deep convolutional neural network model, to convert EEG signals into multidimensional embeddings (). These embeddings are further processed by an MLP classifier to predict object labels (), which correspond to objects present in the image stimuli. The label set matches Imagenet’s labels, available as part of the dataset, representing the most salient object in each image.
The model is trained by minimizing two losses: (A) a categorical cross-entropy loss () between the predicted and ground-truth object labels, and (B) a mean squared error () between the EEG embeddings () and pooled image embeddings () from a pretrained CLIP model Radford et al. (2021), which captures semantically rich image representations. CLIP’s embeddings offer robust transfer learning capabilities, making them ideal for aligning with EEG data.
To better align EEG embeddings with visual stimuli, we simplify images by removing non-central details like color, converting them into sketches using techniques such as Gaussian Blur and Canny filters. While this step is empirical and optional, it helps focus the alignment on core object features.
The overall loss function () balances and , weighted by a hyper-parameter (set to 0.5):
We predict both and for three reasons: (a) aligning EEG embeddings with image embeddings allows us to leverage multimodal vision-language models later, adapting pretrained models for EEG-based text generation, (b) joint optimization ensures the embeddings emphasize salient objects in the images, and (c) object labels, combined with EEG embeddings, can later be fed into multimodal language models to guide more accurate generation. Given the noisy nature of EEG signals, including object labels in the prompts helps keep the model grounded in the salient object and reduces the likelihood of hallucinations.
4.2 Stage2: Priming LLMs with Image Embeddings
To enable LLMs to process multimodal inputs, such as EEG and visual embeddings, we designed a projector inspired by recent advancements in vision-language models Zhang et al. (2024). Since LLMs are inherently text-based and cannot natively accept numerical embeddings, the projector transforms embeddings from the vision and EEG models ( in Stage 1 and in Stage 3) into the token embedding space of the LLM. This ensures the projected embeddings have the same dimensionality as the LLM’s token embeddings.
The projector is a simple feed-forward layer that maps the embeddings into the LLM’s token space. These transformed embeddings are then concatenated with token embeddings extracted from the input prompt, allowing the LLM to process both text and external modality embeddings seamlessly. In our main setting, the input prompt is structured as follows:
In stage 2, we first embed the tokens from the extended prompt, including the object labels, using the LLM’s token embedding layer. The token embeddings, , are then augmented with multimodal embeddings. Specifically, the embedding for the token is replaced by the projected multimodal embedding , and the token embedding for the is replaced by the embedding of the ground-truth object label. The multimodal embedding is computed by projecting external embeddings (e.g., from Stage 1) into the LLM’s token embedding space using the following transformation:
Here, and represent the parameters of the projector. Once is computed, it is prepended to the token embeddings from the input prompt, enabling the LLM to process multimodal information.
During training, labels are created by right-shifting the tokens in the prompt, aligning them with the ground-truth image descriptions. Special tokens such as beginning of sentences (), end of sentences () and padding tokens () are used considering LLM specific tokenizers. The input prompts are also converted into LLM specific chat templates. The LLM is then fine-tuned using a standard cross-entropy loss between the predicted tokens and the actual tokens from the ground-truth descriptions.
4.3 Stage3: Tuning LLMs with EEG Embeddings
This stage closely resembles Stage 2, with the distinction that we utilize , extracted from the EEG encoder trained in Stage 1, to compute the multimodal embeddings . During this stage, the parameters and of the projector undergo further training. We would like to highlight that throughout this and the previous stage, only the projector is trained, while the LLM and EEG encoders remain frozen to mitigate parameter instability caused by EEG noise.
4.4 Inference
During inference, the EEG encoder predicts the object label and EEG embeddings from raw EEG signals. The object labels are concatenated with a generic input prompt to compute token embeddings. The multimodal projector, trained in the previous stages, then computes from . These embeddings are concatenated and fed into LLMs to generate descriptions. Notably, no images are used during inference, making the inference process strictly bimodal.
| ID | Original Image Stimuli | Input | Ground Truth Object | Predicted Object (Stage 1) | Ground Truth Description (GPT-4) | Predicted Description (Thought2Text) |
| 1 |
EEG | mushroom | flower | A large yellow mushroom with a brown stem and a brown cap, surrounded by green foliage. | A group of mushrooms growing on a log. |
| 2 |
EEG + OBJ | piano | piano | A black grand piano in a living room. | A grand piano with a stool in front of it. |
| 3 |
EEG | piano | piano | A man in a red coat and black pants is playing a piano in a room with a chandelier. | A man is playing the piano in a dimly lit room. |
| 4 |
EEG + OBJ | pumpkin | pumpkin | A carved pumpkin with a face and eyes, sitting on a table. | A carved pumpkin with a spooky face on it. |
| 5 |
EEG + OBJ | flower | mushroom | A black and white photograph of a single daisy with a white center and a dark brown center. | A group of mushrooms growing on a log. |
| 6 |
EEG + OBJ | coffee mug | coffee mug | A hand holding a mug with a blue background and a handprint design. | A person holding a coffee mug with the words "World’s Best Dad" written on it. |
| 7 |
EEG | guitar | watch or watches | A young boy sitting on a chair playing a guitar. | A man is holding a guitar in front of a microphone. |
| 8 |
EEG + OBJ | camp or camping | camp or camping | A tent in a mountainous area with trees and fog. | A tent set up in a forest with a campfire nearby. |
5 Experimental Details
5.1 Dataset:
We utilize the open-source CVPR2017 dataset Palazzo et al. (2020) licensed for academic research, featuring EEG signals from six subjects viewing 50 images across 40 ImageNet classes Deng et al. (2009), totaling 2000 images. The data is split into training (), evaluation (), and test () examples. For more details, see Section 3.
5.2 Model Details
We use ChannelNet Palazzo et al. (2020) for the EEG encoder, modifying the final linear layer to produce 512-dimensional embeddings to match the output of the CLIP vision encoder (openai/clip-vit-base-patch32). The EEG encoder is trained with a batch size of 16 for 100 epochs, using the AdamW optimizer and a learning rate of . For fine-tuning the LLM, we use a similar setup: a batch size of 16, training for 5 epochs per stage, and employing gradient accumulation and checkpointing. The learning rate for the LLM fine-tuning is kept at . All implementations are carried out using PyTorch and Huggingface’s transformers library.
We evaluate three LLMs: LLaMa-v3 (meta-llama/Meta-Llama-3-8B-Instruct), Mistral-v03 (mistralai/Mistral-7B-Instruct-v0.3), and Qwen2.57B (Qwen/Qwen2.5-7B-Instruct), selected for their efficiency on consumer-grade GPUs like the NVIDIA RTX 3090. The multimodal embedding is projected to match the token embedding dimensions required by each LLM. All models are permissively licensed for academic research, and training takes approximately 7 GPU hours per LLM training cycle. During inference, we use a batch size of 1, with generation parameters such as top_k, top_p, and temperature set to their default values for each LLM.
To demonstrate the effectiveness of our approach, we compare it against the following baselines: (a) : where LLMs generate a description based solely on the predicted object without any additional input (e.g., if the object is "car," the LLM generates a description of the word "car" following the prompt in Section 4.2); (b) : where we pass a random embedding alongside the predicted object labels to the LLMs; (c) : where the priming step described in Section 4 is skipped; and (d) : where only the EEG embeddings from Stage 1 are used as input, ignoring the object labels. Our proposed Thought2Text solution, which incorporates all stages and all inputs, is labeled ALL in the experiments.
5.3 Evaluation:
We use standard NLG metrics such as BLEU, METEOR, ROUGE Sharma et al. (2017), and BERTScore Zhang et al. (2019). Additionally, we assess generation quality using GPT-4, following Liu et al. (2024b).GPT-4 measures two aspects: fluency, for grammar , and adequacy, for accuracy in conveying meaning, both rated on a scale of 1-5, with 5 denoting highest quality.
6 Results
Table 2 presents a comprehensive comparison of metrics across different models and setups. From these results, it is evident that the complete approach, denoted as ALL, consistently outperforms other setups across all evaluation metrics. As expected, chance-based baselines like and exhibit poor performance. For instance, the LLaMA3-8B_ALL model achieved a BLEU-N (N=1) score of 25.5%, Mistral-7B_ALL scored 26%, and QWEN2.5-7B_ALL reached 22.7%—all significantly higher than their respective chance scores under the chance setups. Notably, some models, particularly the LLaMa variants, show increased sensitivity to random input, leading to further reduction in the scores with chance setups.
We will first discuss the inferences from the results obtained with other alternative setups before delving into a subject-wise analysis.
6.1 Comparison with Stage 2 Omission (NO_STAGE2)
The ROUGE-N (N=1) score for LLaMA3-8B’s variant is 30.0%, whereas the variant achieves only 26.9%, indicating a significant improvement with the inclusion of Stage 2. The same trend is observed for other metrics like BLEU-N (N=1 and N=4), METEOR, and BERT Score, and is observed across other models as well. In terms of adequacy as measured by GPT-4, the variant stands out except for Qwen model which shows a tendency to copy the object and produce shorter sentences in the case of which is positively rated by GPT-4. Overall, the table demonstrates that incorporating Stage 2 (as detailed in Section 4) — which aligns EEG embeddings with image embeddings using a CLIP-based supervision strategy — contributes to higher-quality text generation. The consistently higher BERT scores further emphasize that Stage 2 helps maintains semantic consistency, which aligns with improved Fluency and Adequacy as assessed by GPT-4.
6.2 Generation performance without object labels in the input
While our initial hypothesis was that EEG embeddings—derived from noisy, multichannel EEG data—might not fully capture complex thoughts, thus requiring additional inputs like object labels, it is pleasantly surprising to observe that even without object labels, the models () perform comparably to the best models (). This underscores the effectiveness of aligning EEG embeddings with vision embeddings in Stage 1 and pretraining the LLM, particularly the projectors, in Stage 2 with vision embeddings.
6.3 Subject-wise Analysis
For this analysis, each subject’s EEG data is used to independently train and test the LLMs, simulating a personalized solution. The dataset comprises six subjects, allowing for individual analysis to evaluate the robustness of the approach across different participants.
As depicted in Figure 2, in the subject-wise analysis, the advantages of the complete approach () become even more prominent. Cross-subject and in-subject experiments consistently favor the configuration, with significant improvements in Adequacy score when Stage 2 is included, especially for LLaMA3-8B and Qwen2.5-7B models. These improvements, though numerically small, are crucial in the context of EEG data where every bit of alignment and finetuning matters due to its inherently noisy and sparse nature.
The configuration, which omits the essential alignment step between EEG embeddings and image embeddings, consistently demonstrates lower performance across subjects, as illustrated by the BLEU Unigram scores in Figure 2. This validates our hypothesis that direct fine-tuning without the warm-up provided by Stage 1 and Stage 2 is insufficient for EEG data. The challenge is further exacerbated by the inherent difficulties associated with EEG data, which, even when ethically collected and cleaned, still suffers from limited data availability and significant variability across sessions and subjects.
Subject-wise analysis is essential in practice due to the inherently private and sensitive nature of EEG data. Models like ours, designed for thought-to-text translation, must be developed and deployed within privacy-preserving settings. Creating a personalized solution without access to extensive cross-subject EEG data can be challenging.
To this end, our multi-stage approach demonstrates that pretraining on non-EEG data (such as images) and fine-tuning on small amounts of subject-specific EEG data opens new avenues for privacy-preserving personalized EEG-LLM model development. The efficacy of this approach, as opposed to direct fine-tuning following standard practices, shows that a combination of pretraining and fine-tuning can make personalized EEG models a reality, even with limited data.
6.4 Qualitative Inspection and Basic Error Analysis
For qualitative inspection, we compared the generated descriptions of images produced by both the GPT-4 and the Mistral-7B-v0.3 model ( and variants). Table 3 presents notable examples from our assessment of the model’s generated descriptions. In many instances, the approach with input of EEG + predicted object label generates highly accurate descriptions, as illustrated in positive examples 2 and 4 for the piano and pumpkin, respectively. However, in cases of misgeneration, we encounter not only significantly inaccurate outputs and hallucinations, as seen in examples 6, 7, and 8, but also instances of genuine confusion. For instance, the EEG signal for a flower is misidentified as a mushroom in example 5, leading to incorrect generation. This misidentification indicates potential areas for improvement in object classification.
However, a key advantage of the EEG-only approach is that, even when the predicted object label is incorrect, the model can still produce reasonably coherent descriptions, as seen in example 1. This highlights the robustness of the EEG embeddings in guiding the language model’s generation. Similar observations were made with other models based on LLaMa and QWEN architectures, further validating the consistency and reliability of our approach across various LLM frameworks. We acknowledge that the selected examples are purposefully chosen to illustrate different scenarios.
7 Conclusion and Future Work
In this paper, we introduced a novel approach to convert EEG signals into text, leveraging instruction-tuned LLMs fine-tuned with EEG data. Our method progresses through three stages: training an EEG encoder for feature extraction, fine-tuning LLMs on multimodal data, and further refining them with EEG embeddings for direct text generation from neural signals. Validation on a public EEG dataset demonstrates the efficacy of popular LLMs in translating visual thoughts into text. We observe significant performance improvements compared to chance evaluation, and our methodology, incorporating all stages, performed well in both cross-subject and in-subject analyses, as validated through quantitative evaluation. The qualitative evaluation further provided insights into various scenarios involving EEG signals and object labels versus EEG signals alone as inputs for generating text. These evaluations not only reinforce the effectiveness of our methodology for efficient text generation but also underscore the potential of utilizing EEG data alone to achieve the desired results. However, instances of misidentification that result in incorrect outputs reveal opportunities for further improvements in text generation.
Our future work will focus on optimizing model architectures, leveraging foundational pretrained EEG models on diverse dataset, improving EEG-text alignment through stage 2 training on large scale image datasets and diverse tasks such as optical character recognition, question answering and summarizing and exploring practical applications in healthcare and assistive technologies, marking a significant stride toward accessible "thoughts-to-text" systems.
Limitations
Extracting fine-grained information from EEG signals presents challenges due to high data-to-noise ratio and low spatial resolution. Despite these difficulties, EEG signals can still identify object categories, which can then be used with a generic prompt to aid in text generation. However, misclassification of similar-shaped objects (see the Appendix, Figure 3), such as mistaking a mushroom for a flower, underscores potential ambiguities in object recognition. Additionally, there were instances where the encoder generated unrelated descriptions, such as identifying a coffee maker as a computer or an elephant as a panda. Object classification accuracy varies among subjects (see the Appendix, Table 4) due to individual differences in interpreting images, leading to diverse EEG signal variations and increased prediction variance. One approach to address this is by training personalized models for each subject and assessing their performance. Additionally, implementing methods that enhance the generalizability of predictions across subjects could be explored. Another challenge is data scarcity; Deep Learning models typically require substantial data for training. Hence, acquiring more high-quality, multi-channel EEG data under controlled experimental conditions is crucial to reduce noise. Despite these challenges, our quantitative and qualitative findings demonstrate promising results. We believe that additional training on larger datasets and rigorous controlled experiments will significantly improve performance.
While reading thoughts can be beneficial for individuals with limited ability to communicate, the major risk lies in the potential misuse of BCIs to intrude into thoughts without consent. However, with appropriate measures and regulations, these risks should not hinder advancements in understanding and translating human cognition, as the benefits outweigh the challenges.
Ethics Statement
For this work, we utilized anonymized open-source EEG data, acknowledging the sensitivity of EEG data and the imperative of ethical compliance. All experimental data used in our research were anonymized to protect participant privacy and uphold ethical standards.